### v1

@@ 2.2 Intrinsic time delays in the gamma-ray range
Solving Eq. 5 gives access to the evolution of lepton spectra with time, to the evolution of spectral energy distributions as well as to the associated multi-wavelength light curves, after transformation from the source frame to the observer frame. An example of such distributions can be found in Fig. 1 for a typical weak flare based on a pure one-zone SSC emission scenario, which can serve as a good proxy for stationary states and VHE flares of BL Lacs. A standard extragalactic background light (EBL) model provided by Kneiske et al. (2002, 2004) is taken into account for all the work presented in this paper.
Figure 2: Intrinsic time delays at gamma-ray energies obtained for a pure SSC scenario, for the set of parameters given in Table 1 (black curve, 87 mG) and for two other values of the magnetic field, 50 mG (blue) and 130 mG (red). The different trends - increasing in blue, flat in black and decreasing in red - are here obtained only from the variation of the magnetic field amplitude ${B}||$ | |, and start to appear at $E\sim 100$ MeV.
@@
Figure 3: Upper panel: intrinsic time delay evolution for adiabatic expansion and compression effects, introduced with the variation of the blob’s radius $R$. Lower panel: illustration of delay evolution for decreasing external photon field effects, modelled as a blackbody emitted by an accretion disc and further reprocessed by a BLR, varied by the increasing distance $d$ between the central black hole and the emitting blob. For the SSC part, the simulations are based on the set of parameters given in Table 1. For both plots, the reference time for lag calculation is taken at 1 MeV (see the main text for further explanations).
Each light curve can be characterised by a typical arrival time of photons, chosen to be the time at which the light curve peaks. The evolution of light curves over energy thus gives access to the evolution of the arrival times over energy. To estimate temporal delays as per energy band centred on energy $E$, we introduce a reference arrival time noted $t_{\mathrm{ref}}$. The time delays are then given by
@@
We are primarily interested in the time delay evolution with respect to energy. These intrinsic delays navigate between two types of regimes at high energy ($E>100$ MeV) in the pure SSC scenario: increasing and decreasing trends in the TeV range. These two regimes, respectively called long-lasting acceleration (or acceleration-driven), and fast acceleration (or cooling-driven) regimes in Perennes et al. (2020), arise from an imbalance between acceleration and cooling processes. Leptons start to emit photons as soon as they get accelerated such that both processes take effect simultaneously but with different relative influence. A transition zone between the two regimes, here referred to as ”flat trend”, can arise from the combination and overall balance between the two processes. In this transition zone, intrinsic delays do not evolve and seem to cancel out leading to a flat curve showing no variation with the photon energy. An example of intrinsic time delay evolution with energy can be found in Fig. 2 where $t_{\mathrm{ref}}=t_{E=1\mathrm{MeV}}$. The three trends are represented, here obtained by varying the initial value of the magnetic field.
@@
In this work, and in order to investigate the impact of adiabatic and EIC processes on the intrinsic delays, we define a set of reference parameters permissible for a typical TeV-emitting blazar, with the intention to reproduce the special case of a flat trend (Table 1). This choice reflects the fact that no significant time delay at TeV energies in AGNs has been detected yet, with the notable exception of the 4$\pm$1 minutes delay detected by MAGIC in 2005 for the archetypal BL Lac object Mkn 501 at redshift $z=0.03$ (Albert et al., 2007). The flat regime therefore appears as a better fit to observations. This state is fairly unstable and slight variations of parameters can send delays along an increasing or decreasing trend, and can thus be used to better highlight the impact of individual model parameters on intrinsic delays.
@@
The impact of adiabatic processes is illustrated in Fig. 3 (upper panel) which shows the variation of intrinsic time delays with respect to the blob’s radius evolution index $m_{r}$. The reference value $m_{r}=0$ (the radius stays constant) is represented by the black curve. Red points correspond to expansion of the blob ($m_{r}>0$), while blue points correspond to a compression ($m_{r}<0$). When the blob expands with time, as expected in the absence of an efficient confinement mechanism, its plasma is diluted and leptons diffuse their energy. This leads to a non-radiative additional cooling effect placing delays in the long lasting acceleration regime with an increasing trend. On the contrary, if the blob is compressed during the flare and its volume decreases with time, its plasma gets more condensed and is heated up. This rise in energy translates into a gain in kinetic energy, and leptons are naturally accelerated, placing delays in the fast acceleration regime with a decreasing trend. Then, it appears adiabatic processes do not modify the overall description of intrinsic time delays and only influence the regime in which the delays evolve with energy for a specific given flare.
@@
The evolution of intrinsic time delays with respect to an external photon field contribution is shown in Fig. 3 (lower panel). All the parameters governing the external field are kept constant with the exception of the blob-accretion disc distance $d$. In this configuration with non zero EIC, numerical computations can diverge when the adiabatic expansion index $m_{r}$ is smaller or equal to 1 (Levy, 2021). For that reason, we adopt the value $m_{r}=1.001$. The other parameters are set to the values given in Table 1. At very large distances $d$, the case $d=\infty$ is equivalent to a vanishing EIC contribution and the lag evolution follows an increasing trend due to the adiabatic expansion effect. In our description, reducing the distance $d$ corresponds to increasing the energy density of the external photon field and the importance of the EIC radiation and cooling (see Eq. 10). The presence of additional EIC simply increases the radiative cooling, without modifying the acceleration term. This induces a global evolution of the intrinsic delays and pushes them towards the decreasing trend at VHE. The situation is a bit more complex at intermediate distances ($d=15\times 10^{17}$ cm) and intermediate gamma-ray energies where delays first increase for $E<1$ GeV, decrease between $E=1$ GeV and $E=1$ TeV, and increase again for $E>1$ TeV. This effect results from the combination of the three different cooling mechanisms (SSC, EIC, and adiabatic expansion), which are more or less important depending on the energy, relatively to the acceleration process. However, such behaviour does not call into question the presence of the two main regimes already identified, with increasing or decreasing trends, which characterise the intrinsic time delays for general leptonic scenarios at TeV energies and above, where LIV effects could be expected.
@@
We can conclude that the presence of adiabatic processes and of a significant EIC emission do not modify the overall picture of intrinsic leptonic time delays in the gamma-ray range, so FSRQs should follow similar intrinsic patterns as BL Lacs. However it can significantly modify the regime of the intrinsic delays. Significant EIC ultimately pushes intrinsic delays towards the decreasing trend (cooling-driven regime) at VHE, and adiabatic compression does the same, while adiabatic expansion pushes intrinsic delays towards the increasing trend (long-lasting acceleration regime). The model described here yields intrinsic delays quite effortlessly from the variation of source physical quantities and properties. A variation of about 50% in the value of the main model parameters can indeed be enough to induce intrinsic time delays larger than 200 s/TeV (as shown in Fig. 2 for the magnetic field), which could be potentially detectable, depending on the experimental errors. At a first glance, this might appear in contradiction with the lack of detected time delays at TeV energies blazar data collected so far. However, such a non-detection is usually explained by a lack of precision of the current measurements, so that the uncertainty on time lag measurements are still too large, leading to values compatible with zero. This question should be directly answered by the next generation of Cherenkov instruments such as the Cherenkov Telescope Array (Rosales De Leon et al., 2024). If no time delay were ever detected at VHE gamma-rays, this would impose strong constraints on standard leptonic flare models and could put them in difficulty and even discard some of them.
@@
3 Exploring multi-wavelength leptonic flares with an additional LIV delay
3.1 Injection of LIV delays
@@
where $E_{\mathrm{LC}}$ is the mean light curve energy, and $\tau$ is the linear LIV term (Eq. 3). It can be emphasised that this transformation does not change the shape of individual light curves, but it modifies total time delays as well as the temporal evolution of the SEDs which therefore have to be rebuilt from the light curves in order to include the LIV effect. $\tau$ is kept as a free parameter expressed in s TeV-1 and can take positive (subluminal LIV effect) or negative (superluminal LIV effect) values. A typical uncertainty on measured parameter $\tau$ equivalent to a $1\sigma$ confidence level for Mkn 501 is $\sim 200$ s TeV-1 (Bolmont et al., 2022, Table A1) so it was chosen as a typical value for the injected lag.
With such considerations, total time delays observed for a flare with an intrinsic increasing trend and for various LIV contributions can be found in Fig. 4. Intrinsic effects are shown in black, while intrinsic combined with LIV effects are shown in red for $\tau>0$, and blue for $\tau<0$. As LIV contribution depends on energy, having $\tau>0$ tends to send delays towards the increasing trend while $\tau<0$ send them towards the decreasing one. Therefore, LIV either amplifies time delays when both effects impose the same evolution with energy, or tends to suppress time delays when the two effects impose opposite evolution. For $\tau$ big enough, LIV effects can even change the observed trend at high energies. Then, it appears that separating LIV-induced and intrinsic delays can be challenging without a specific strategy.
Figure 4: Simultaneous description of intrinsic and LIV effects in time delays at high energies, illustrated by varying the importance of LIV around the case (in black) of the intrinsic increasing trend from Fig. 2 with $B_{0}=50$ mG. Positive (superluminal) and negative (subluminal) values of $\tau$ are represented in red and blue respectively.
@@ 3.2 Multi-wavelength X-ray and $\gamma$-ray flares from blazars
The study of intrinsic time delays performed in Sect. 2 focuses on the gamma-ray domain. This limitation was motivated by the fact LIV effects can only become significant at very high energies. However, intrinsic delays also arise from the low energy emission as discussed for instance by Lewis et al. (2016). Indeed, within leptonic scenarios, the inverse Compton high-energy bump is directly dependent on the synchrotron low-energy one. A strong correlation between the two bumps of the SEDs is expected since it is the same population of high energy leptons which produces the synchrotron emission, and the high energy emission due to the inverse Compton process. The correlation is even greater for pure SSC models where the ambient scattered photons are the synchrotron photons emitted themselves by the same high energy leptons. In any case, the two bumps are intimately related and evolve together in a similar fashion for leptonic scenarios. This similarity between energy bands should then also appear in light curves and more importantly in intrinsic time delays. Since LIV effects should only appear in the high energy domain, studying the relationship between sets of intrinsic delays arising in low and high energy ranges could provide strategies based on multi-wavelength observations to discriminate between intrinsic and LIV-induced time delays. The occurrence of LIV effects should result in breaking or weakening the similarity between observed low-energy and high-energy time delays.
@@
To test this hypothesis, we extended the energy range to derive intrinsic delays at X-ray, optical and infrared energies. To better separate low and high energy domains, a special care was given to the choice of the reference arrival time $t_{\mathrm{ref}}=t_{E=E_{\mathrm{ref}}}$ corresponding to a specific energy $E_{\mathrm{ref}}$ (Eq. 11). The energy of the dip in SEDs, the low flux region between the two bumps, was chosen for this quantity. Since the location of the dip varies with the time evolution of the generated SEDs, the SED with the highest $\nu F_{\nu}$ flux value at that location was considered as a reference. The energy $E_{\mathrm{ref}}$ is then taken as the energy of the dip for that specific SED. $E_{\mathrm{ref}}$ typically varies between 0.01 and 1 MeV for BL Lacs. To highlight the relationship between the two energy domains, the parameter space and the different regimes of delays were explored. One illustrative example is shown in Fig. 5 for a typical flare of a BL Lac object at low and high redshifts. A clear similarity between the two energy domains (synchrotron and inverse Compton) appears in the absence of any LIV effect. This behaviour was found to be a systematic one: both sets of low and high energy delays always follow the same trend whatever the set of parameters used to generate them. Such systematic similarity indicates the Klein-Nishina and EBL effects which tend to modify the observed fluxes at the highest energies, do not significantly affect the intrinsic delays, and by extension their regime of evolution with energy. In the case of FSRQs, the symmetry between synchrotron and inverse-Compton domains remains but is weakened compared to the pure SSC case due to the introduction of an external photon field. The external inverse-Compton emission then gain an additional component independent from the synchrotron one, which somewhat reduces the inter-dependency between the delays in the X-ray and gamma-ray domains.
@@
At high redshift $(z=1)$, observed fluxes become much smaller for the same set of source parameters, and the EBL contribution is much stronger hence suppressing the very high energy tail of the SEDs. In order to be compatible with CTA sensitivity (Mazin, 2019), a selection was made where a threshold is applied to the observable flux. Only data sets with flux $F>10^{-20}\ \mathrm{erg\,cm^{-2}\,s^{-1}}$ are used in our analysis. As a consequence, modifying the redshift changes the selection on the data sets as can be seen in Fig. 5 where the time delay evolution (increasing regime) with redshift $z=1$ (orange points) is compared to the one with the reference redshift $z=0.03$ (black points). While a large redshift simply eliminates information at VHE and thus removes the data points, the regime and shape of the distributions are preserved. The values of the delays change only slightly both in the X-ray and gamma-ray domains. The same qualitative results hold independently of the regime of the intrinsic time delays. This illustrates the weak dependency between intrinsic time delays and redshift.
@@
Figure 5: Intrinsic time delays extended to low energy domains for the case of the intrinsic increasing trend from Fig. 2, obtained for small (z = 0.03, black) and large redshift (z = 1, orange). Here the energy reference value is taken at $E_{\mathrm{ref}}=1$ MeV. At large redshift, fluxes at the highest energies are reduced and corresponding data points suppressed due to strong EBL absorption. The regime found at low redshift is retained nonetheless. No LIV delay is injected here.

### v2

@@ 2.2 Intrinsic time delays in the gamma-ray range
Solving Eq. 5 gives access to the evolution of lepton spectra with time, to the evolution of spectral energy distributions, and to the associated multi-wavelength light curves, after transformation from the source frame to the observer frame. Examples of these distributions can be found in Fig. 1 for a typical weak flare based on a pure one-zone SSC emission scenario, which can serve as a good proxy for stationary states and VHE flares of BL Lacs. A standard extragalactic background light (EBL) model provided by Kneiske et al. (2002, 2004) is taken into account for all the work presented in this paper.
Figure 2: Intrinsic time delays at gamma-ray energies obtained for a pure SSC scenario, for the set of parameters given in Table 1 (black curve, 87 mG) and for two other values of the magnetic field, 50 mG (blue) and 130 mG (red). The different trends (increasing in blue, flat in black, and decreasing in red) are here obtained only from the variation of the magnetic field amplitude ${B}||$ | |, and start to appear at $E\sim 100$ MeV.
@@
Figure 3: Intrinsic time delay as a function of the energy. Upper panel: Intrinsic time delay evolution for adiabatic expansion and compression effects, introduced with the variation of the blob’s radius $R$. Lower panel: Illustration of delay evolution for decreasing external photon field effects, modelled as a blackbody emitted by an accretion disc, and further reprocessed by a BLR, varied by the increasing distance $d$ between the central black hole and the emitting blob. For the SSC part, the simulations are based on the set of parameters given in Table 1. For both plots, the reference time for lag calculation is taken at 1 MeV (see the main text for further explanations).
Each light curve can be characterised by a typical arrival time of photons, chosen to be the time at which the light curve peaks. The evolution of light curves over energy thus gives access to the evolution of the arrival times over energy. To estimate temporal delays as per energy band centred on energy $E$, we introduce a reference arrival time denoted $t_{\mathrm{ref}}$. The time delays are then given by
@@
We are primarily interested in the time delay evolution with respect to energy. These intrinsic delays navigate between two types of regimes at high energy ($E>100$ MeV) in the pure SSC scenario: increasing and decreasing trends in the TeV range. These two regimes, respectively called long-lasting acceleration (or acceleration-driven), and fast acceleration (or cooling-driven) regimes in Perennes et al. (2020), arise from an imbalance between acceleration and cooling processes. Leptons start to emit photons as soon as they are accelerated, such that the two processes take effect simultaneously, but with different relative influence. A transition zone between the two regimes, here referred to as flat trend, can arise from the combination and overall balance between the two processes. In this transition zone, intrinsic delays do not evolve and seem to cancel out, leading to a flat curve showing no variation with the photon energy. An example of intrinsic time delay evolution with energy can be found in Fig. 2, where $t_{\mathrm{ref}}=t_{E=1\mathrm{MeV}}$. The three trends are represented, here obtained by varying the initial value of the magnetic field.
@@
In this work, and in order to investigate the impact of adiabatic and EIC processes on the intrinsic delays, we define a set of reference parameters permissible for a typical TeV-emitting blazar, with the intention of reproducing the special case of a flat trend (Table 1). This choice reflects the fact that no significant time delay at TeV energies in AGN has been detected yet, with the notable exception of the 4$\pm$1 min delay detected by MAGIC in 2005 for the archetypal BL Lac object Mkn 501 at redshift $z=0.03$ (Albert et al., 2007). The flat regime therefore appears to be a better fit to observations. This state is fairly unstable and slight variations of parameters can send delays along an increasing or decreasing trend, and can thus be used to better highlight the impact of individual model parameters on intrinsic delays.
@@
The impact of adiabatic processes is illustrated in Fig. 3 (upper panel), which shows the variation of intrinsic time delays with respect to the blob’s radius evolution index $m_{r}$. The reference value $m_{r}=0$ (the radius stays constant) is represented by the black curve. The red points correspond to expansion of the blob ($m_{r}>0$), while the blue points correspond to a compression ($m_{r}<0$). When the blob expands with time, as expected in the absence of an efficient confinement mechanism, its plasma is diluted and leptons diffuse their energy. This leads to a non-radiative additional cooling effect placing delays in the long lasting acceleration regime with an increasing trend. On the contrary, if the blob is compressed during the flare and its volume decreases with time, its plasma becomes more condensed and is heated up. This rise in energy translates into a gain in kinetic energy, and leptons are naturally accelerated, placing delays in the fast acceleration regime with a decreasing trend. Thus, it appears adiabatic processes do not modify the overall description of intrinsic time delays and only influence the regime in which the delays evolve with energy for a specific given flare.
@@
The evolution of intrinsic time delays with respect to an external photon field contribution is shown in Fig. 3 (lower panel). All the parameters governing the external field are kept constant with the exception of the blob-accretion disc distance $d$. In this configuration with non-zero EIC, numerical computations can diverge when the adiabatic expansion index $m_{r}$ is smaller than or equal to 1 (Levy, 2021). For that reason, we adopt the value $m_{r}=1.001$. The other parameters are set to the values given in Table 1. At very large distances $d$, the case $d=\infty$ is equivalent to a vanishing EIC contribution and the lag evolution follows an increasing trend due to the adiabatic expansion effect. In our description, reducing the distance $d$ corresponds to increasing the energy density of the external photon field and the importance of the EIC radiation and cooling (see Eq. 10). The presence of additional EIC simply increases the radiative cooling without modifying the acceleration term. This induces a global evolution of the intrinsic delays and pushes them towards the decreasing trend at VHE. The situation is a bit more complex at intermediate distances ($d=15\times 10^{17}$ cm) and intermediate gamma-ray energies where delays first increase for $E<1$ GeV, decrease between $E=1$ GeV and $E=1$ TeV, and increase again for $E>1$ TeV. This effect results from the combination of the three different cooling mechanisms (SSC, EIC, and adiabatic expansion), which are more or less important depending on the energy, relative to the acceleration process. However, this behaviour does not call into question the presence of the two main regimes already identified, with increasing or decreasing trends, which characterise the intrinsic time delays for general leptonic scenarios at TeV energies and above, where LIV effects could be expected.
@@
We can conclude that the presence of adiabatic processes and of a significant EIC emission do not modify the overall picture of intrinsic leptonic time delays in the gamma-ray range, so FSRQs should follow intrinsic patterns similar to those of BL Lacs. However, it can significantly modify the regime of the intrinsic delays. Significant EIC ultimately pushes intrinsic delays towards the decreasing trend (cooling-driven regime) at VHE, and adiabatic compression does the same, while adiabatic expansion pushes intrinsic delays towards the increasing trend (long-lasting acceleration regime). The model described here yields intrinsic delays quite effortlessly from the variation in source physical quantities and properties. A variation of about 50% in the value of the main model parameters can indeed be enough to induce intrinsic time delays larger than 200 s/TeV (as shown in Fig. 2 for the magnetic field), which could be potentially detectable, depending on the experimental errors. At a first glance, this might appear to be in contradiction with the lack of detected time delays at TeV energies blazar data collected so far. However, such a non-detection is usually explained by a lack of precision of the current measurements, so that the uncertainty on time lag measurements is still too large, leading to values compatible with zero. This question should be directly answered by the next generation of Cherenkov instruments, such as the Cherenkov Telescope Array (Rosales De Leon et al., 2024). If no time delay were ever detected at VHE gamma rays, this would impose strong constraints on standard leptonic flare models and could put them in difficulty and even discard some of them.
@@
3 Exploring multi-wavelength leptonic flares with an additional Lorentz invariance violation delay
3.1 Injection of Lorentz invariance violation delays
@@
where $E_{\mathrm{LC}}$ is the mean light curve energy and $\tau$ is the linear LIV term (Eq. 3). It should be noted that this transformation does not change the shape of individual light curves, but it modifies total time delays as well as the temporal evolution of the SEDs, which therefore have to be rebuilt from the light curves in order to include the LIV effect. The parameter $\tau$ is kept as a free parameter expressed in s TeV-1, and it can take positive (subluminal LIV effect) or negative (superluminal LIV effect) values. A typical uncertainty on measured parameter $\tau$ equivalent to a $1\sigma$ confidence level for Mkn 501 is $\sim 200$ s TeV-1 (Bolmont et al., 2022, Table A1), and so it was chosen as a typical value for the injected lag.
With such considerations, total time delays observed for a flare with an intrinsic increasing trend and for various LIV contributions can be found in Fig. 4. Intrinsic effects are shown in black, while intrinsic combined with LIV effects are shown in red for $\tau>0$ and blue for $\tau<0$. As LIV contribution depends on energy, having $\tau>0$ tends to send delays towards the increasing trend, while $\tau<0$ sends them towards the decreasing trend. Therefore, LIV either amplifies time delays when both effects impose the same evolution with energy, or tends to suppress time delays when the two effects impose opposite evolution. For $\tau$ that is large enough, LIV effects can even change the observed trend at high energies. Then, it appears that separating LIV-induced and intrinsic delays can be challenging without a specific strategy.
Figure 4: Simultaneous description of intrinsic and LIV effects in time delays at high energies, illustrated by varying the importance of LIV around the case of the intrinsic increasing trend from Fig. 2 with $B_{0}=50$ mG (in black). Positive (superluminal) and negative (subluminal) values of $\tau$ are represented in red and blue, respectively.
@@ 3.2 Multi-wavelength X-ray and $\gamma$-ray flares from blazars
The study of intrinsic time delays performed in Sect. 2 focuses on the gamma-ray domain. This limitation was motivated by the fact LIV effects can only become significant at very high energies. However, intrinsic delays also arise from the low-energy emission as discussed for instance by Lewis et al. (2016). Within leptonic scenarios, the inverse-Compton high-energy bump is directly dependent on the synchrotron low-energy bump. A strong correlation between the two bumps of the SEDs is expected since it is the same population of high-energy leptons that produces the synchrotron emission and the high-energy emission due to the inverse-Compton process. The correlation is even greater for pure SSC models where the ambient scattered photons are the synchrotron photons emitted themselves by the same high-energy leptons. In any case, the two bumps are closely related and evolve together in a similar fashion for leptonic scenarios. This similarity between energy bands should then also appear in light curves and more importantly in intrinsic time delays. Since LIV effects should only appear in the high-energy domain, studying the relationship between sets of intrinsic delays arising in low- and high-energy ranges could provide strategies based on multi-wavelength observations to discriminate between intrinsic and LIV-induced time delays. The occurrence of LIV effects should result in breaking or weakening the similarity between observed low-energy and high-energy time delays.
@@
To test this hypothesis, we extended the energy range to derive intrinsic delays at X-ray, optical, and infrared energies. To better separate low- and high-energy domains, special care was given to the choice of the reference arrival time $t_{\mathrm{ref}}=t_{E=E_{\mathrm{ref}}}$ corresponding to a specific energy $E_{\mathrm{ref}}$ (Eq. 11). The energy of the dip in SEDs, the low-flux region between the two bumps, was chosen for this quantity. Since the location of the dip varies with the time evolution of the generated SEDs, the SED with the highest $\nu F_{\nu}$ flux value at that location was considered as a reference. The energy $E_{\mathrm{ref}}$ is then taken as the energy of the dip for that specific SED. The value of $E_{\mathrm{ref}}$ typically varies between 0.01 and 1 MeV for BL Lacs. To highlight the relationship between the two energy domains, the parameter space and the different regimes of delays were explored. One illustrative example is shown in Fig. 5 for a typical flare of a BL Lac object at low and high redshifts. A clear similarity between the two energy domains (synchrotron and inverse-Compton) appears in the absence of any LIV effects. This behaviour was found to be systematic: both sets of low- and high-energy delays always follow the same trend whatever the set of parameters used to generate them. This systematic similarity indicates that the Klein-Nishina and EBL effects, which tend to modify the observed fluxes at the highest energies, do not significantly affect the intrinsic delays, and by extension their regime of evolution with energy. In the case of FSRQs, the symmetry between synchrotron and inverse-Compton domains remains, but is weakened compared to the pure SSC case, due to the introduction of an external photon field. The external inverse-Compton emission then gains an additional component independent from the synchrotron component, which somewhat reduces the inter-dependency between the delays in the X-ray and gamma-ray domains.
@@
At high redshift $(z=1)$, observed fluxes become much smaller for the same set of source parameters, and the EBL contribution is much stronger, and hence suppresses the very high-energy tail of the SEDs. In order to be compatible with CTA sensitivity (Mazin, 2019), a selection was made where a threshold is applied to the observable flux. Only datasets with flux $F>10^{-20}\ \mathrm{erg\,cm^{-2}\,s^{-1}}$ are used in our analysis. As a consequence, modifying the redshift changes the selection on the datasets, as can be seen in Fig. 5 where the time delay evolution (increasing regime) with redshift $z=1$ (orange points) is compared to that with the reference redshift $z=0.03$ (black points). While a high redshift simply eliminates information at VHE, and thus removes the data points, the regime and shape of the distributions are preserved. The values of the delays change only slightly in the X-ray and the gamma-ray domains. The same qualitative results hold independently of the regime of the intrinsic time delays. This illustrates the weak dependency between intrinsic time delays and redshift.
@@
Figure 5: Intrinsic time delays extended to low-energy domains for the case of the intrinsic increasing trend from Fig. 2, obtained for low redshift (z = 0.03, black) and high redshift (z = 1, orange). Here the energy reference value is taken at $E_{\mathrm{ref}}=1$ MeV. At high redshift, fluxes at the highest energies are reduced and the corresponding data points are suppressed due to strong EBL absorption. The regime found at low redshift is retained nonetheless. No LIV delay is injected here.